according to newton's law of cooling the rate at which a substance cools in moving air is proportional to the difference between the temperature of the substance and that of the year if the temperature of the year is 298 k a and the substance cools from 478 k to 428 k in 20 minutes. find when the temperature will be 405 k

1
Expert's answer
2022-03-11T13:42:10-0500

Solution;

According to Newton's law of cooling;

"-\frac{dT}{dt}\alpha(T-T_s)"

"\frac{dT}{dt}=-k(T-T_s)"

Integrating;

"\int\frac{dT}{T-T_s}=-k\int dt"

"ln(T-T_s)=-kt"

"T-T_s=e^{-kt}+C"

"T(t)=T_s+Ce^{-kt}"

"T_s=298k"

Boundary conditions;

"t=0,T=478k"

"t=20,T=428"

"478=298+Ce^0"

"C=478-298=180"

Also;

"\frac{478}{428}=\frac{Ce^o}{Ce^{-20k}}"

"e^{20k}=1.116"

"k=0.00552"

Therefore;

"T(t)=298+180e^{-0.00552t}"

At 405,t=?;

"405=298+180e^{-0.00552t}"

"107=180e^{-0.00552t}"

"ln(\frac{107}{180})=-0.0055t"

"t=94.23minutes"